OpenStudy (anonymous):

Help please this is my last year and i really need to do good! I will fan and medal!!! Given f(x)=2x^2+8 and g(x)=5x-6, find each function for questions 4-5. 4. (f+g)(x) 5. (fxg)(x)

OpenStudy (breyliance):

i would suggest contacting your teacher

OpenStudy (anonymous):

I cant he is out of the office today

OpenStudy (breyliance):

ask a parent interweb idk

OpenStudy (anonymous):

\[(f+g)(x)=f(x) + g(x)\] Add the functions and combine like terms \[f(x) + g(x)=2x^2+8+5x-6\]

OpenStudy (breyliance):

holy crap

OpenStudy (anonymous):

:( none of my family knows this stuff and either do i, i have trouble understanding the problems

OpenStudy (anonymous):

5. \((f \times g)(x)=f(x) \times g(x)=(2x^2+8)(5x-6)\) Use FOIL/Distribution to simplify

OpenStudy (skullpatrol):